Assume 100.00 g of the compound what is the mass of each element in 100.00 g of compound what are the moles of each element in 100.00 g of compound.

1
Expert's answer
2021-12-08T01:43:02-0500

Since the identity of the substance in question is not given, I will use a substance such as Sodium carbonate, Na2CO3 to demonstrate how the question can be answered.

Using RAM of Na=23, C=12, and O=16)

RFM/MM of Na2CO3 = (23x2) + 12 +(16x3) = 109g/mol


Mass of each element in 100g of the compound;


Mass of Na = (46/109) x 100g = 42.2g

Mass of C = (12/109) x 100g = 11.0g

Mass of O = (48/109) x 100g = 44.0g


Moles of each element in the 100g of the compound;


Moles of Na = 42.2/23 = 1.83mol

Moles of C = 11.0/12 = 0.92mol

Moles of O = 44.0/16 = 2.75mol